OpenStudy (anonymous):

A 0.50 kilogram ball is held at a height of 20 meters. What is the kinetic energy of the ball when it reaches halfway after being released?

OpenStudy (anonymous):

initial velocity u=0 h=10 m g=9.8 m/s final velocity v=? \[v^2-u^2=2gh\] find v

OpenStudy (anonymous):

then Kinetic energy\[=\frac{ 1 }{ 2 } m v^2\]